MAT 145: Topics In Contemporary Math
19: Solving Exponential Growth

4) The population of bison was almost hunted to extinction when the United States
expanded westward. The population decreased by 40% per year when the expansion
began in 1803. If the herds of bison originally contained 500,000 bison, long would it
take for there to only be 1,000 bison left?

Answer by ankor@dixie-net.com(22740)

The population of bison was almost hunted to extinction when the United States
expanded westward.
The population decreased by 40% per year when the expansion began in 1803.
If the herds of bison originally contained 500,000 bison, long would it
take for there to only be 1,000 bison left?
:
let t = time required for this to happen
With 40% decrease, that means .6 of original amt remain
:
500000*.6^t = 1000
.6^t = 1000/500000
.6^t = 1/500
uing natural logs
t*ln(.6) = ln(1/500)
t = ln(1/500)/ln(.6)
using calc
t ~ 12 years